# Digital Simulation of Non-Abelian Parafermions in Superconducting Circuits

### v4

@@ III.1 Ground state preparation
$H^{(d)}\left|{j}\right\rangle=\frac{1}{\sqrt{3}}\sum_{i=1}^{d-1}\omega^{ij}\left|{i}\right\rangle.$ (5)
@@
After applying the $H^{(3)}$ to each representative qutrit, all the light plaquettes (including the light boundary plaquettes) are prepared in the ground state [Fig. 2(c)]. Next, we select the control qutrit on each dark bulk plaquette [green dots in Fig. 2(b)] and apply $CZ$, $CZ^{\dagger}$, and $CX$ gates to prepare all the dark plaquettes in the ground state [Figs. 2(d) and 2(e)]. The controlled two-qutrit gates must be applied layer by layer, ensuring that the state of the representative qutrit remains unchanged until the controlled operations on their respective plaquettes are applied. Ground state preparation on the dark plaquettes does not cause any light plaquette to deviate from the ground state, as their stabilizers commute with each other. In this way, we prepare the entire system in the ground state. The depth of the ground state preparation circuit depends linearly on the size of the lattice ($\sqrt{n}$ in our set up, with $n$ the number of qutrits).
@@
where tilde indicates that the conditional gate $\tilde{C}O_{P}$ is controlled by the ancillary qutrit. The gate $\tilde{C}O_{P}$ consists of $\tilde{C}Z_{1}$, $\tilde{C}X_{2}$, $\tilde{C}Z_{3}^{\dagger}$, and $\tilde{C}X_{4}^{\dagger}$ gates, where the subscripts denote the order of qutrits in a stabilizer, as defined in Eq. (1). After applying a $H^{(3)}$ gate to each ancillary qutrit, we measure all the ancillary states at the same time to project each of them onto $\left|{\tilde{0}}\right\rangle$. The plaquette state transforms as
@@
where $\tilde{H}^{(3)}$ indicates that the operator acts on the ancillary qutrit, and $\tilde{P}(0)$ is a projector that projects the ancillary qutrit onto the $\left|{\tilde{0}}\right\rangle$ state. Therefore, after the procedure described in Eqs. (7), (8), and (9), each plaquette is subjected to a ground-state projector and prepared in the ground state. This method can be directly generalized to the more general qudit cases.
@@ III.2 Creation of parafermions
An alternative method arises in the presence of natural holes in the lattice [Fig. 2(g)], which result from imperfections in the geometric arrangement of qutrits due to fabrication constraints in chip manufacturing. Exploiting this imperfection, a pair of parafermions can be created by preparing the ground state on the imperfect lattice. However, for the region containing the hole, additional controlled gates, such as the controlled-$XZ$ gate, are required to enforce the ground-state condition. An example featuring a minimal hole is illustrated in Fig. 2(g).
@@ IV FUSION OPERATIONs OF THE GENERATED PARAFERMIONS
In this section, we propose the scheme to realize the fusion of parafermions. The previous section presents a method for creating parafermion pairs through engineering dislocations. Nevertheless, since the parafermions remain adjacent, studying the fusion rules of an individual parafermion is difficult because the nearby parafermion can influence the fusion and introduce ambiguity in the interpretation of results. To address this, we generalize the code deformation method to the qudit setting, deriving the operators that facilitate the spatial movement of parafermions. This generalization is technically non-trivial due to the non-Hermitian nature of the stabilizers, requiring careful treatment of both the stabilizer and its Hermitian conjugate. Physically, this property is intrinsically linked to the non-self-duality of anyons and the nonlinearity of parafermion operator transformations, a connection we leave for future work. The generalized code deformation method yields operators that enable the horizontal and vertical movement of parafermions along the lattice. This capability serves as a powerful tool for parafermion manipulation and further facilitates essential operations, such as braiding [see Sec. V]. Given its significance and fundamental role in this study, we first introduce the generalized code deformation method before combining it with string operators to establish a protocol for realizing the fusion of parafermions.
@@ IV.1 Generalized code deformation
Code deformation is a method used to convert one error-correcting code into another by modifying the stabilizers, enabling fault-tolerant logical gate operations within surface codes [92, 93]. Beyond quantum information science, code deformation also finds applications in condensed matter physics. Topological surface codes, such as the toric code, are rooted in topological quantum field theory, which provides the framework for topological orders. In this context, code deformation involves altering the topological structure of the field theory, such as engineering topological defects, to enrich the properties of topological orders [94]. While the method has been extensively studied in qubit systems, its application to qudit systems remains relatively unexplored. In this subsection, we generalize the code deformation method to the qudit system. This generalization enables the derivation of unitary operators that facilitate the spatial movement of parafermions across the lattice. Our work extends the applicability of code deformation to a broader class of qudit surface codes.
@@
Figure 3: Generalized code deformation and the fusion of parafermions. (a) Code deformation that moves the parafermion vertically. (b) Code deformation that moves the parafermion horizontally. (c) Illustration of the creation of a composite excitation and braiding $e$ or $\overline{e}$ with one parafermion. After the braiding, the excitations fuse to the trivial topological charge. (d) Explicit procedure for realizing the fusion of parafermions. Observables are characterized by $\langle\Pi_{p^{\prime}}^{d-1}\rangle=\langle\Pi_{p}^{1}\rangle=1$. (e) Quantum circuit and two-qutrit gates used in the procedure.
@@
Before proceeding to the realization of the operators $U_{h}$ and $U_{v}$ using quantum circuits, we briefly review the code deformation method used to move Majorana modes in stabilizer codes with SC qubits [3, 4] and compare it with our generalization. In that case, both the original stabilizer $S_{\text{old}}$ and the transformed stabilizer $S_{\text{new}}$ after moving the Majorana mode are Hermitian, simplifying the construction of a unitary operator that transforms the system into an eigenvalue-$1$ eigenstate of the new stabilizer. Specifically, since $S_{\text{old}}$ anticommutes with $S_{\text{new}}$, their product $iS_{\text{new}}S_{\text{old}}$ is Hermitian, motivating the ansatz of the unitary operator $e^{\theta S_{\text{new}}S_{\text{old}}}$, where the parameter $\theta=\frac{\pi}{4}$ is fixed by applying the operator to the eigenvalue-$1$ eigenstate of $S_{\text{old}}$ and requiring that the resulting state be the eigenvalue-$1$ eigenstate of $S_{\text{new}}$. Nevertheless, in the qudit case, this ansatz does not apply. The absence of Hermitian stabilizers complicates the construction of the unitary operator in Eq. (10), making the generalization significantly non-trivial.
@@
where $\theta_{1}=\frac{\pi}{4}$ and $\theta_{2}\approx 1.696\pi$. These rotation and diagonal matrices can be experimentally implemented using microwave drives for level transitions and phase shifts of drive pulses [82, 83]. The circuit implementation for $U_{v}$ is not necessarily unique and optimization algorithms can enhance its efficiency [95]. Machine learning methods may offer further potential improvements. Once $U_{v}$ is implemented, the operator $U_{h}$ can be obtained using the circuit in Fig. 4(b), where the qutrit SWAP gate is implemented using the circuit in Fig. 4(c). Here, the $K^{(3)}$ gate is defined by $K^{(3)}\left|{j}\right\rangle=\left|{3-j}\right\rangle$.
@@ IV.2 Realization of fusion operations
In Fig. 3(d), we explicitly illustrate the procedure for realizing the fusions in Eq. (17). The protocol begins by measuring two horizontally adjacent qutrits in the $XZ^{\dagger}$ basis and projecting them onto the eigenvalue-$1$ eigenstate, thereby creating a pair of parafermions. Since the parafermions are created from the ground state, their initial state is in a fixed fusion channel: $\left|{\gamma,\gamma;1}\right\rangle$. We then separate the two parafermions by moving one downward using the operator $U_{v}$. Next, an $e\overline{m}$ dyon excitation is created by applying two Pauli operators. The $e$ excitation is then braided around one parafermion using a string operator, which consists of a sequence of Pauli operators. After braiding, the $e$ excitation is transformed into the $m$ excitation and proceeds to annihilate the $\overline{m}$ excitation. Once this process is completed, we move the parafermion back to a neighboring position and measure the expectation values $\langle\Pi^{d-1}_{p^{\prime}}\rangle$ and $\langle\Pi^{1}_{p}\rangle$ by the QST method [the (vi) step in Fig. 3(d)]. Before fusion, both $\langle\Pi^{d-1}_{p^{\prime}}\rangle$ and $\langle\Pi^{1}_{p}\rangle$ vanish due to the initial fusion channel $\left|{\gamma,\gamma;1}\right\rangle$. After fusion, the parafermon absorbs the $e\overline{m}$ excitation, resulting in the predicted expectation values $\overline{\langle\Pi^{d-1}_{p^{\prime}}\rangle}=1$ and $\overline{\langle\Pi^{1}_{p}\rangle}=1$, averaged over multiple experimental realizations. A similar procedure can be applied to create and braid the $\overline{e}m$ excitation, providing an experimentally accessible observable for realizing the fusion rules in Eq. (17). Notably, reversing steps (iii-iv) in Fig. 3(d) implements the fusion rule in Eq. (16). In Fig. 3(e), we demonstrate the corresponding circuit and two-qutrit operators used in the procedure.
@@ V Braiding of parafermion modes
To encode information, a minimum number of four parafermions is required [Fig. 5(a)]. In this configuration, there are only two parity operators, $\Lambda_{1}=\overline{\omega}\gamma_{1}\gamma_{2}^{\dagger}$ and $\Lambda_{3}=\overline{\omega}\gamma_{3}\gamma_{4}^{\dagger}$ with $\gamma$ the parafermion operator [see Appendix. B]. For convenience, the computational space is restricted to the two-dimensional subspace spanned by $\Lambda_{1}\Lambda_{3}=1$. With this restriction, four parafermions are sufficient to encode a logical qutrit. To find the basis of the computational space, we diagonalize the parity operator $\Lambda_{i}$ and obtain the eigenstates $\{\left|{0}\right\rangle_{\Lambda_{i}},\left|{1}\right\rangle_{\Lambda_{i}},\left|{2}\right\rangle_{\Lambda_{i}}\}$. The logical qutrit state is defined by:
@@
On the other hand, physically, the basis of the computational space is defined by the fusion tree of the four parafermions [Fig. 5(b)]. The specific transformation between the basis defined by the eigenstates of the parity operators and the fusion basis $\left|{\gamma_{1},\gamma_{2};1/e\overline{m}/\overline{e}m}\right\rangle\otimes\left|{\gamma_{3},\gamma_{4};1/\overline{e}m/e\overline{m}}\right\rangle$ requires selecting two base points on the lattice and constructing the parafermion operators using string operators whose start and end points are at these base points and satisfying the algebra in Eq. (24[18, 3]. Nevertheless, in this work, we do not make this exact transformation by selecting base points. Instead, we take a more direct approach to measure the braiding consequences of the parafermions, which does not rely on the explicit relation between the fusion basis and the logical qutrit basis $\left|{j}\right\rangle_{\Lambda_{1}}\otimes\left|{3-j}\right\rangle_{\Lambda_{3}}$.
In the logical qutrit basis $\{\left|{\overline{j}}\right\rangle\}$, the half-braiding operator $U_{2}$ that exchanges $\gamma_{2}$ and $\gamma_{3}$ is an off-diagonal matrix, exhibiting the non-Abelian statistics. In a specific representation, the matrix elements of $U_{2}$ are given by: $\langle{\overline{i}}|U_{2}\left|{\overline{j}}\right\rangle=\frac{1}{\sqrt{3}}c_{i-j},$ where $c_{j}=\omega^{\frac{j(j+5)}{2}}$ [see Appendix. B]. The full braiding matrix in the logical qutrit basis is:
@@
We now move on to the explicit procedure. In Fig. 5(d), we show the implementation of full braiding of parafermions. The process begins by creating four parafermion $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$, and $\gamma_{4}$ from the ground state, arranged in the configuration shown in Fig. 5(a). In the ground state, measuring $\langle\Pi_{p_{i}}^{0}\rangle$ in these plaquettes yields $\langle\Pi_{p_{i}}^{0}\rangle=1$. We then move $\gamma_{3}$ to the bottom, followed by moving $\gamma_{2}$ to the right. After these steps, we move $\gamma_{2}$ back to its original position and finally return $\gamma_{3}$ to its starting point. After braiding, the parafermions $\gamma_{1}$ and $\gamma_{2}$ fuse to form a composite excitation, denoted by $f$, which consists of the components $1$, $\overline{e}m$, and $e\overline{m}$. Similarly, parafermions $\gamma_{3}$ and $\gamma_{4}$ fuse into another composite excitation $f^{\prime}$ [Fig. 5(c)]. In general, $f$ and $f^{\prime}$ are distinct, but they are related by an electric-magnetic transformation, where $e\leftrightarrow m$. According to Eq. (20), the probability of measuring the topologically trivial charge $1$ from the composite excitation $f$ or $f^{\prime}$ is always $\frac{1}{3}$. Therefore, when measuring the expectation value $\langle\Pi_{p_{i}}^{0}\rangle$ using the QST method on the plaquettes that support the dislocation, the averaged result over multiple repeated experiments is given by $\overline{\langle\Pi_{p_{i}}^{0}\rangle}=\frac{1}{3}$. This provides an experimental observable for the braiding of parafermions. In Fig. 5(e), we show the two-qutrit gates used in this procedure.

### v5

@@ III.1 Ground state preparation
$H^{(d)}\left|{j}\right\rangle=\frac{1}{\sqrt{d}}\sum_{i=0}^{d-1}\omega^{ij}\left|{i}\right\rangle.$ (5)
@@
After applying the $H^{(3)}$ to each representative qutrit, all the light plaquettes (including the light boundary plaquettes) are prepared in the ground state [Fig. 2(c)]. Next, we select the control qutrit on each dark bulk plaquette [green dots in Fig. 2(b)] and apply $CZ^{\dagger}$, $CZ$, and $CX$ gates to prepare all the dark plaquettes in the ground state [Figs. 2(d) and 2(e)]. The controlled two-qutrit gates must be applied layer by layer, ensuring that the state of the representative qutrit remains unchanged until the controlled operations on their respective plaquettes are applied. Ground state preparation on the dark plaquettes does not cause any light plaquette to deviate from the ground state, as their stabilizers commute with each other. In this way, we prepare the entire system in the ground state. The depth of the ground state preparation circuit depends linearly on the size of the lattice ($\sqrt{n}$ in our set up, with $n$ the number of qutrits).
@@
where tilde indicates that the conditional gate $\tilde{C}O_{P}$ is controlled by the ancillary qutrit. The gate $\tilde{C}O_{P}$ consists of $\tilde{C}Z_{1}$, $\tilde{C}X_{2}$, $\tilde{C}Z_{3}^{\dagger}$, and $\tilde{C}X_{4}^{\dagger}$ gates, where the subscripts denote the order of qutrits in a stabilizer, as defined in Eq. (1). After applying a generalized Hadamard gate to each ancillary qutrit, we measure all ancillary states at the same time to project each of them onto $\left|{\tilde{0}}\right\rangle$. The plaquette state transforms as
@@
where $\tilde{H}^{(3)}$ indicates that the generalized Hadamard gate acts on the ancillary qutrit, and $\tilde{P}(0)$ is a projector that projects the ancillary qutrit onto the $\left|{\tilde{0}}\right\rangle$ state. Therefore, after the procedure described in Eqs. (7), (8), and (9), each plaquette is subjected to a ground-state projector and prepared in the ground state. This method can be directly generalized to the more general qudit cases.
@@ III.2 Creation of parafermions
A method arises when natural holes appear in the lattice [Fig. 2(g)], potentially resulting from imperfections in the geometric arrangement of qutrits due to fabrication constraints. Exploiting this imperfection, a pair of parafermions can be created by preparing the ground state on the imperfect lattice. For the surface containing the hole, however, additional controlled gates, such as the controlled-$XZ$ gate, are required to enforce the ground-state condition. An example featuring a minimal hole is illustrated in Fig. 2(g).
Figure 3: Generalized code deformation and the fusion of parafermions. (a) Code deformation that moves the parafermion vertically. (b) Code deformation that moves the parafermion horizontally. (c) Illustration of the creation of a composite excitation and braiding an $e$ or $\overline{e}$ excitation with one parafermion. After the braiding, the excitations fuse to the trivial topological charge. (d) Explicit procedure for realizing the fusion of parafermions. Observables are characterized by $\langle\Pi_{p^{\prime}}^{d-1}\rangle=\langle\Pi_{p}^{1}\rangle=1$. (e) Quantum circuit and two-qutrit gates used in the procedure.
@@ IV FUSION OPERATIONs OF THE GENERATED PARAFERMIONS
In this section, we propose a scheme to realize parafermion fusion. While the previous section presented a method to create parafermion pairs via engineered dislocations, the adjacent creation of parafermions hinders the study of fusion rules of individual parafermion, as nearby parafermion can influence the fusion result. To address this, we generalize the code deformation method to qudit systems and derive operators that enable spatial separation of parafermions. This generalization is technically nontrivial due to the non-Hermitian nature of the stabilizers, requiring careful treatment of both the stabilizer and its Hermitian conjugate. Physically, this complexity is intrinsically linked to the non-self-duality of anyons and the nonlinearity of parafermion operator transformations—a connection we leave to future work. The resulting code deformation protocol allows both horizontal and vertical movement of parafermions on the surface, offering a versatile tool for their manipulation and enabling further operations such as braiding (see Sec. V). Given its significance and fundamental role in this study, we first introduce the generalized code deformation method before combining it with string operators to establish a protocol for realizing the fusion of parafermions.
@@ IV.1 Generalized code deformation
Code deformation is a method used to convert one error-correcting code into another by modifying the stabilizers, enabling fault-tolerant logical gate operations of surface codes [152, 153]. Beyond quantum information science, code deformation also finds applications in condensed matter physics. Topological surface codes, such as the toric code, are rooted in topological quantum field theory, which provides the general framework of topological orders. In this context, code deformation involves altering the topological structure of the field theory, such as engineering topological defects, to enrich the properties of topological orders [133]. While the method has been extensively studied in qubit systems, its application to qudit systems remains relatively unexplored. In this subsection, we generalize the code deformation method to the qudit system. This generalization enables the derivation of unitary operators that facilitate the spatial movement of parafermions on the surface. Our work extends the applicability of code deformation to a broader class of qudit surface codes.
@@
Before proceeding to the realization of the operators $U_{h}$ and $U_{v}$ using quantum circuits, we briefly review the code deformation method used to move Majorana modes in stabilizer codes with SC qubits [42, 43] and compare it with our generalization. In the qubit case, both the original and deformed stabilizers $S_{\text{old}}$ and $S_{\text{new}}$ are Hermitian. This allows a straightforward construction of a unitary operator that maps an eigenstate of $S_{\text{old}}$ with eigenvalue $+1$ to an eigenstate of $S_{\text{new}}$ with the same eigenvalue. Specifically, since $S_{\text{old}}$ and $S_{\text{new}}$ anticommute, the product $iS_{\text{new}}S_{\text{old}}$ is Hermitian, motivating the construction of $e^{\frac{\pi}{4}S_{\text{new}}S_{\text{old}}}$. Applying this operator achieves the desired state transformation. In the qudit case, however, the stabilizers are no longer Hermitian, and this strategy breaks down. The lack of Hermiticity introduces significant complications in constructing the unitary operator defined in Eq. (10), making the generalization highly non-trivial.
@@
where $\theta_{1}=\frac{\pi}{4}$ and $\theta_{2}\approx 1.696\pi$. These rotation and diagonal matrices can be experimentally implemented using microwave drives for level transitions and phase shifts of drive pulses [144, 145]. The circuit implementation for $U_{v}$ is not necessarily unique and optimization algorithms can enhance its efficiency [24]. Machine learning methods may offer further improvements. Once $U_{v}$ is implemented, the operator $U_{h}$ can be obtained using the circuit in Fig. 4(b), where the qutrit SWAP gate is implemented using the circuit in Fig. 4(c). Here, the $K^{(3)}$ gate is defined by $K^{(3)}\left|{j}\right\rangle=\left|{3-j}\right\rangle$.
@@ IV.2 Realization of fusion operations
In Fig. 3(d), we explicitly illustrate the procedure to perform the fusions in Eq. (17). The protocol begins by measuring two horizontally adjacent qutrits in the $XZ^{\dagger}$ basis and projecting them onto the eigenvalue-$1$ eigenstate, thereby creating a pair of parafermions. Since parafermions are created from the ground state, their initial state is in a fixed fusion channel: $\left|{\gamma,\gamma;1}\right\rangle$. We then separate the two parafermions by moving one downward using the operator $U_{v}$. Next, an $e\overline{m}$ dyon excitation is created by applying two Pauli operators. The $e$ excitation is then braided around one parafermion using a string operator, which consists of a sequence of Pauli operators. After braiding, the $e$ excitation is transformed into the $m$ excitation and proceeds to annihilate the $\overline{m}$ excitation. Once this process is completed, we move the parafermion back to a neighboring position and measure the expectation values $\langle\Pi^{d-1}_{p^{\prime}}\rangle$ and $\langle\Pi^{1}_{p}\rangle$ by the QST method [the (vi) step in Fig. 3(d)]. Before fusion, both $\langle\Pi^{d-1}_{p^{\prime}}\rangle$ and $\langle\Pi^{1}_{p}\rangle$ vanish due to the initial fusion channel $\left|{\gamma,\gamma;1}\right\rangle$. After fusion, the parafermon absorbs the $e\overline{m}$ excitation, resulting in the predicted expectation values $\overline{\langle\Pi^{d-1}_{p^{\prime}}\rangle}=1$ and $\overline{\langle\Pi^{1}_{p}\rangle}=1$, averaged over multiple experimental realizations. A similar procedure can be applied to create and braid the $\overline{e}m$ excitation, providing an experimentally accessible observable to realize the fusion rules in Eq. (17). In particular, reversing steps (iii-iv) in Fig. 3(d) implements the fusion rule in Eq. (16). In Fig. 3(e), we demonstrate the corresponding circuit and two-qutrit operators used in the procedure.
@@ V Braiding of parafermion modes
To encode information, a minimum number of four parafermions is required [Fig. 5(a)]. In this configuration, there are only two parity operators, $\Lambda_{1}=\overline{\omega}\gamma_{1}\gamma_{2}^{\dagger}$ and $\Lambda_{3}=\overline{\omega}\gamma_{3}\gamma_{4}^{\dagger}$ with $\gamma$ the parafermion operator (see Appendix. B). For convenience, the computational space is restricted to the two-dimensional subspace spanned by $\Lambda_{1}\Lambda_{3}=1$. With this restriction, four parafermions are sufficient to encode a logical qutrit. To find the basis of the computational space, we diagonalize the parity operator $\Lambda_{i}$ and obtain the eigenstates $\{\left|{0}\right\rangle_{\Lambda_{i}},\left|{1}\right\rangle_{\Lambda_{i}},\left|{2}\right\rangle_{\Lambda_{i}}\}$. The logical qutrit state is defined by:
@@
On the other hand, physically, the basis of the computational space is defined by the fusion tree of the four parafermions [Fig. 5(b)]. The specific transformation between the basis defined by the eigenstates of the parity operators and the fusion basis $\left|{\gamma_{1},\gamma_{2};1/e\overline{m}/\overline{e}m}\right\rangle\otimes\left|{\gamma_{3},\gamma_{4};1/\overline{e}m/e\overline{m}}\right\rangle$ involves selecting two base points on the surface and constructing the parafermion operators using string operators connecting these base points, with carefully selected arrows to satisfy commutation relations Eq. (24[40, 42]. Notably, the representation of parafermion operators via string operators is reminiscent of the Jordan-Wigner transformation, highlighting their non-locality. This non-local nature stems from the defining non-local commutation relations of parafermions Eq. (24). Since these string operators share common endpoints, the parity operators Eq. (25) function as loop operators, measuring the fusion results of parafermions enclosed within the loop. Consequently, the logical qutrit is globally defined when the parafermions are sufficiently separated. This is also evident in the fusion operations discussed in Sec. IV.2, where transforming the logical qutrit state involves creating a composite excitation $e\overline{m}$ adjacent to the boundary and braiding either $e$ or $\overline{m}$ around the parafermion. When parafermions are engineered at the center of the system, such fusion operations require a string operator connecting the boundary to the parafermion defect and a loop operator encircling the defect. These operations consist of a set of local operators whose number, i.e., the code distance, increases linearly with system size, thereby ensuring topological protection of the logical qutrit state. In this study, we adopt a more direct approach to measure the braiding statistics of parafermions without relying on the explicit construction of parity operators and the detailed relationship between the fusion basis and the logical qutrit basis.
In the logical qutrit basis $\{\left|{\overline{j}}\right\rangle\}$, the half-braiding operator $U_{2}$ that exchanges $\gamma_{2}$ and $\gamma_{3}$ is an off-diagonal matrix, exhibiting the non-Abelian statistics. In a specific representation, the matrix elements of $U_{2}$ are given by: $\langle{\overline{i}}|U_{2}\left|{\overline{j}}\right\rangle=\frac{1}{\sqrt{3}}c_{i-j},$ where $c_{j}=\omega^{\frac{j(j+5)}{2}}$ (see Appendix. B). The full braiding matrix in the logical qutrit basis is:
@@
We now move on to the explicit procedure. In Fig. 5(d), we show the implementation of full braiding of parafermions. The process begins by creating four parafermion $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$, and $\gamma_{4}$ from the ground state, arranged in the configuration shown in Fig. 5(a). In the ground state, measuring $\langle\Pi_{p_{i}}^{0}\rangle$ in these plaquettes yields $\langle\Pi_{p_{i}}^{0}\rangle=1$. We then move $\gamma_{3}$ to the bottom, followed by moving $\gamma_{2}$ to the right. After these steps, we move $\gamma_{2}$ back to its original position and finally return $\gamma_{3}$ to its starting point. After braiding, the parafermions $\gamma_{1}$ and $\gamma_{2}$ fuse to form a composite excitation, denoted by $f$, which consists of the components $1$, $\overline{e}m$, and $e\overline{m}$. Similarly, parafermions $\gamma_{3}$ and $\gamma_{4}$ fuse into another composite excitation $f^{\prime}$ [Fig. 5(c)]. In general, $f$ and $f^{\prime}$ are distinct, but they are related by an electric-magnetic transformation, where $e\leftrightarrow m$. According to Eq. (20), the probability of measuring the topologically trivial charge $1$ from the composite excitation $f$ or $f^{\prime}$ is always $\frac{1}{3}$. Therefore, when measuring the expectation value $\langle\Pi_{p_{i}}^{0}\rangle$ using the QST method on the plaquettes that support the dislocation, the averaged result over multiple repeated experiments is given by $\overline{\langle\Pi_{p_{i}}^{0}\rangle}=\frac{1}{3}$. This provides an experimental observable for the braiding of parafermions. In Fig. 5(e), we show two-qutrit gates used in the procedure.